# xGR: Efficient Generative Recommendation Serving at Scale

### v1

@@ 1 Introduction
Challenge 2: how to maintain beam search efficiency while minimizing both sorting and filtering overhead in high-concurrency recommendation scenarios? The core operation of beam search is to select beam width (BWBW) sequences from a large pool of candidate sequences [meister2020if]. For each identified token, only the Top-$K$ likely candidate tokens for the next step are retained [leblond2021machine]. Accurate recommendations require larger BWBW and $K$ values, but the sorting and filtering operations become particularly time-consuming. As the decode phase progresses, beam search continuously generates new candidate sequences and discards old sequences. Frequent creation and destruction of related data structures incurs significant overhead. Moreover, due to the numerous possibilities of token ID combinations, not all of them correspond to actual products, leading to candidate sequence invalidation [kong2025minionerec].
@@ 2.2.2 Beam Search Algorithm
Beam search is a heuristic search algorithm widely adopted in GR to balance the trade-off between retrieval quality and computational cost. It maintains a set of BWBW most promising partial sequences, where BWBW is the beam width.
As shown in Figure 2, at each time step $t$, the algorithm expands all BWBW active beams by computing the probability distribution over the item vocabulary. To manage the search space, the system first selects the Top-$K$ most likely next-token candidates for each beam. From this aggregated pool of $BW\times K$ candidates, it then identifies the global Top-BWBW sequences with the highest cumulative log-probabilities to form the new set of active beams for step $t+1$. This procedure repeats until the generated item identifiers are complete. Unlike typical LLM tasks where BWBW is small, generative recommendation demands high diversity, often requiring large values for both BWBW and $K$ (e.g., $128\times 128$ or even $512\times 512$). This results in a massive candidate pool (up to $2.6\times 10^{5}$ candidates per time step), transforming the lightweight sorting operation into a computationally heavy bottleneck.
@@ 5 Attention Computation
Attention computation remains the primary performance bottleneck in GR scenarios. Specifically, the process is divided into one prefill phase and three decode phases, with each decode phase generating a token ID (TIDTID), where the resulting TIDTID triplet represents an item ID. Unlike prior study that identifies the prefill phase as the dominant cost [du2025prefillonly], large BWBW and $K$ settings make the decode stage also particularly time-consuming. In the following, we present the KV cache management and task allocation mechanisms.
@@ 5.1 KV Cache Management
Figure 7 illustrates the KV cache management throughout the whole inference process. The shared cache stores the token-level outputs of the prompt produced during the prefill phase, whereas the unshared cache holds the BWBW token IDs generated at each decode phase. Specifically, since the number of decode phases (NDND) is known in advance, xAttention initializes the unshared cache size to exactly the product of BWBW and NDND, thereby avoiding redundant block loading. In addition, the unshared cache is also managed at the token granularity without extra alignment, avoiding the overhead of block copying. After each decode phase, the beam prefixes are updated and extended with new token IDs.
@@ 5.2 Staged Computation Allocation
Given that accelerators typically contain dozens of CGs, effectively partitioning CGs across stages to achieve load balance is critical. On the one hand, the computational volume of the shared stage is generally greater than that of the unshared stage, potentially requiring more CGs. On the other hand, the shared cache only needs to be loaded once, whereas the unshared cache must be updated at each decode step. The limited CGs allocated to the unshared stage exacerbate intra-CG cache conflicts and complicate the scratchpad memory management. As for the merging stage, the optimal CG allocation is strongly correlated with the preceding stages and cannot be predetermined. To this end, xAttention employs a lightweight decision tree regressor [chen2016xgboost] to predict the performance of each CG partition setting represented as triplets. In addition to the partition setting, the input parameters also include the lengths of unshared and shared caches, i.e., the number of tokens. Note that in mature deployments, GR parameters such as BWBW and $K$, as well as attention parameters like head size, are fixed. Therefore, there is no need to include them as inputs to the regressor. Consequently, the overall cost of collecting the training set and model training is feasible.
@@ 6 Beam Search
In GR scenarios, beam search is essential for personalized recommendations to users from a vast item space. In each decode phase, each beam selects token IDs with Top-$K$ probabilities from the vocabulary to expand its sequence. Then, all beam sequences are traversed to obtain the tokens with Top-BWBW probabilities to enter the next decode phase. However, naive beam search implementation is highly time-consuming and easily violates SLO requirements. To address this issue, we introduce the following optimization techniques.
@@ 6.2 Early Sorting Termination
During token generation, beam search accumulates log-probabilities (log_prob) rather than multiplying raw probabilities to ensure numerical stability. After that, beam search selects the Top-BWBW token sequences from the filtered candidate sequences. This procedure can be viewed as a partial sorting problem, which obviates the need for fully sorting and thereby provides an opportunity to reduce computational complexity. Another feature is that the log_prob results for each beam are inherently in descending order. Based on the above considerations, xBeam adopts an early termination mechanism to improve sorting efficiency.
@@
Figure 11 shows the sorting process with early termination. xBeam maintains a global min heap of size BWBW to store the top-ranking token sequences along with their associated log_prob values. During sorting, xBeam visits the leaves of each sub-beam tree sequentially. During the traversal of each beam, if the leaf’s log_prob exceeds that of the heap’s top element, it is inserted into the heap, and the heap structure is adjusted to maintain order. Otherwise, the sorting operation of that beam is terminated immediately. After traversing all beams, xBeam retrieves the top-BWBW token sequences. In this way, xBeam significantly reduces sorting overhead while ensuring high-quality sequence selection.
@@ 6.3 Data Structure Reuse
Beam search typically requires a larger BWBW (e.g., 512) to ensure accurate recommendations. However, continuously generating new candidate sequences while discarding old ones leads to substantial waste of computational and memory resources. This is because the frequent creation and destruction of sequence-related data structures, as well as the associated data movement operations, incur significant overhead. Fortunately, since the BWBW is predetermined and fixed throughout the beam search process, this provides an opportunity for resource reuse. Specifically, during the generation of new candidate sequences, xBeam does not allocate entirely new data structure; instead, it reuses the data structure previously occupied by old sequences. Moreover, once beam search completes, xBeam updates the memory regions corresponding to the old sequences with the content of the newly generated ones. This strategy not only reduces memory consumption but also lowers the resource management overhead.
@@ 9.3 Memory Efficiency
We further evaluate memory scalability under varying input lengths (Figure 16). With the BWBW fixed at 256, xGR maintains a remarkably low footprint—peaking at only 12.0GB even with 3k tokens—significantly outperforming xLLM, which consistently consumes around 30GB. This confirms that xGR decouples memory usage from sequence length, ensuring robustness for long-context recommendation tasks.

### v2

@@ 1 Introduction
Challenge 2: how to maintain beam search efficiency while minimizing both sorting and filtering overhead in high-concurrency recommendation scenarios? The core operation of beam search is to select beam width ($BW$) sequences from a large pool of candidate sequences [meister2020if]. For each identified token, only the Top-$K$ likely candidate tokens for the next step are retained [leblond2021machine]. Accurate recommendations require larger $BW$ and $K$ values, but the sorting and filtering operations become particularly time-consuming. As the decode phase progresses, beam search continuously generates new candidate sequences and discards old sequences. Frequent creation and destruction of related data structures incurs significant overhead. Moreover, due to the numerous possibilities of token ID combinations, not all of them correspond to actual products, leading to candidate sequence invalidation [kong2025minionerec].
@@ 2.2.2 Beam Search Algorithm
Beam search is a heuristic search algorithm widely adopted in GR to balance the trade-off between retrieval quality and computational cost. It maintains a set of $BW$ most promising partial sequences, where $BW$ is the beam width.
As shown in Figure 2, at each time step $t$, the algorithm expands all $BW$ active beams by computing the probability distribution over the item vocabulary. To manage the search space, the system first selects the Top-$K$ most likely next-token candidates for each beam. From this aggregated pool of $BW\times K$ candidates, it then identifies the global Top-$BW$ sequences with the highest cumulative log-probabilities to form the new set of active beams for step $t+1$. This procedure repeats until the generated item identifiers are complete. Unlike typical LLM tasks where $BW$ is small, generative recommendation demands high diversity, often requiring large values for both $BW$ and $K$ (e.g., $128\times 128$ or even $512\times 512$). This results in a massive candidate pool (up to $2.6\times 10^{5}$ candidates per time step), transforming the lightweight sorting operation into a computationally heavy bottleneck.
@@ 5 Attention Computation
Attention computation remains the primary performance bottleneck in GR scenarios. Specifically, the process is divided into one prefill phase and three decode phases, with each decode phase generating a token ID ($TID$), where the resulting $TID$ triplet represents an item ID. Unlike prior study that identifies the prefill phase as the dominant cost [du2025prefillonly], large $BW$ and $K$ settings make the decode stage also particularly time-consuming. In the following, we present the KV cache management and task allocation mechanisms.
@@ 5.1 KV Cache Management
Figure 7 illustrates the KV cache management throughout the whole inference process. The shared cache stores the token-level outputs of the prompt produced during the prefill phase, whereas the unshared cache holds the $BW$ token IDs generated at each decode phase. Specifically, since the number of decode phases ($ND$) is known in advance, xAttention initializes the unshared cache size to exactly the product of $BW$ and $ND$, thereby avoiding redundant block loading. In addition, the unshared cache is also managed at the token granularity without extra alignment, avoiding the overhead of block copying. After each decode phase, the beam prefixes are updated and extended with new token IDs.
@@ 5.2 Staged Computation Allocation
Given that accelerators typically contain dozens of CGs, effectively partitioning CGs across stages to achieve load balance is critical. On the one hand, the computational volume of the shared stage is generally greater than that of the unshared stage, potentially requiring more CGs. On the other hand, the shared cache only needs to be loaded once, whereas the unshared cache must be updated at each decode step. The limited CGs allocated to the unshared stage exacerbate intra-CG cache conflicts and complicate the scratchpad memory management. As for the merging stage, the optimal CG allocation is strongly correlated with the preceding stages and cannot be predetermined. To this end, xAttention employs a lightweight decision tree regressor [chen2016xgboost] to predict the performance of each CG partition setting represented as triplets. In addition to the partition setting, the input parameters also include the lengths of unshared and shared caches, i.e., the number of tokens. Note that in mature deployments, GR parameters such as $BW$ and $K$, as well as attention parameters like head size, are fixed. Therefore, there is no need to include them as inputs to the regressor. Consequently, the overall cost of collecting the training set and model training is feasible.
@@ 6 Beam Search
In GR scenarios, beam search is essential for personalized recommendations to users from a vast item space. In each decode phase, each beam selects token IDs with Top-$K$ probabilities from the vocabulary to expand its sequence. Then, all beam sequences are traversed to obtain the tokens with Top-$BW$ probabilities to enter the next decode phase. However, naive beam search implementation is highly time-consuming and easily violates SLO requirements. To address this issue, we introduce the following optimization techniques.
@@ 6.2 Early Sorting Termination
During token generation, beam search accumulates log-probabilities (log_prob) rather than multiplying raw probabilities to ensure numerical stability. After that, beam search selects the Top-$BW$ token sequences from the filtered candidate sequences. This procedure can be viewed as a partial sorting problem, which obviates the need for fully sorting and thereby provides an opportunity to reduce computational complexity. Another feature is that the log_prob results for each beam are inherently in descending order. Based on the above considerations, xBeam adopts an early termination mechanism to improve sorting efficiency.
@@
Figure 11 shows the sorting process with early termination. xBeam maintains a global min heap of size $BW$ to store the top-ranking token sequences along with their associated log_prob values. During sorting, xBeam visits the leaves of each sub-beam tree sequentially. During the traversal of each beam, if the leaf’s log_prob exceeds that of the heap’s top element, it is inserted into the heap, and the heap structure is adjusted to maintain order. Otherwise, the sorting operation of that beam is terminated immediately. After traversing all beams, xBeam retrieves the top-$BW$ token sequences. In this way, xBeam significantly reduces sorting overhead while ensuring high-quality sequence selection.
@@ 6.3 Data Structure Reuse
Beam search typically requires a larger $BW$ (e.g., 512) to ensure accurate recommendations. However, continuously generating new candidate sequences while discarding old ones leads to substantial waste of computational and memory resources. This is because the frequent creation and destruction of sequence-related data structures, as well as the associated data movement operations, incur significant overhead. Fortunately, since the $BW$ is predetermined and fixed throughout the beam search process, this provides an opportunity for resource reuse. Specifically, during the generation of new candidate sequences, xBeam does not allocate entirely new data structure; instead, it reuses the data structure previously occupied by old sequences. Moreover, once beam search completes, xBeam updates the memory regions corresponding to the old sequences with the content of the newly generated ones. This strategy not only reduces memory consumption but also lowers the resource management overhead.
@@ 9.3 Memory Efficiency
We further evaluate memory scalability under varying input lengths (Figure 16). With the $BW$ fixed at 256, xGR maintains a remarkably low footprint—peaking at only 12.0GB even with 3k tokens—significantly outperforming xLLM, which consistently consumes around 30GB. This confirms that xGR decouples memory usage from sequence length, ensuring robustness for long-context recommendation tasks.